OpenStudy (anonymous):

If f(x) = x2 - 1, find f(3). 5 2 8 -2

OpenStudy (anonymous):

f(3)=3^2-1=9-1=8

OpenStudy (anonymous):

so its 8? @marsss

OpenStudy (anonymous):

yes it's

OpenStudy (anonymous):

thank you :) @marsss

OpenStudy (anonymous):

you're welcome baby :)